OpenStudy (anonymous):

simplify sin(30+x)+cos(60+x) I can't find any identity that is related to this?

OpenStudy (usukidoll):

ummm looks familiar though

OpenStudy (usukidoll):

sin (A + B) = (sin A)(cos B) + (cos A)(sin B) sin (A - B) = (sin A)(cos B) - (cos A)(sin B) tan (A + B) = (tan A + tan B)/(1 - (tan A)(tan B)) tan (A - B) = (tan A - tan B)/(1 + (tan A)(tan B))

OpenStudy (anonymous):

first find X then plug into a scientific calculator: sin(30+_) + cos(60+_), then you get you answer!

OpenStudy (anonymous):

how do I find x?

OpenStudy (jdoe0001):

http://www.pleacher.com/mp/mgifs/gifs1/sumdif.gif

OpenStudy (usukidoll):

let's split this up for a bit...

OpenStudy (jdoe0001):

well, your will be in terms of "x"

OpenStudy (jdoe0001):

well, your will be in terms of "x" ^ answer =)

OpenStudy (anonymous):

most likely by grouping like terms, so x=_+_

OpenStudy (usukidoll):

sin(30+x)+cos(60+x) since we have a + in sin sin (A + B) = (sin A)(cos B) + (cos A)(sin B) so a = 30 and b = x MOreover... we have a + in cos cosacosb-sinasinb where a = 60 and b = x

OpenStudy (usukidoll):

sin(30+x)+cos(60+x) sin(30+x) =sin30cosx+cos30sinx cos(60+x) = cos60cosx+sin60sinx

OpenStudy (usukidoll):

sorry I'm having trouble reading this... *zooms in*

OpenStudy (usukidoll):

sin30cosx+cos30sinx + cos60cosx+sin60sinx o_O

OpenStudy (anonymous):

how does that help though? doesnt it make it more confusin? or do I convert sin30 into the trig ratios?

OpenStudy (usukidoll):